Circular No. 7958

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002eo IN NGC 710 W. D. Li, University of California at Berkeley, reports the LOTOSS (cf.

IAUC 7906) discovery of an apparent supernova (mag ∼ 18.0) on unfiltered KAIT CCD images taken on Aug. 20.5 and 21.5 UT. SN 2002eo is located at α = 1^{h}52^{m}55^{s}.18, δ = +36^{o}03′22′′.2 (equinox 2000.0), which is 13′′.9 east and 8′′.8 north of the nucleus of NGC 710. A KAIT image taken on Aug. 14.4 showed nothing at this position (limiting mag ∼ 19.0).

V838 MONOCEROTIS A. Henden, Universities Space Research Association and U.S. Naval

Observatory (USNO); and U. Munari, Asiago Astronomical Observatory, Padua, report that radical changes now affect the light echo they discovered around V838 Mon (cf. IAUC 7859). New R-band imaging, obtained on Aug. 22.5 UT with the USNO Flagstaff Station 1.0-m telescope (scale 0′′.68/pixel) with the object now emerging from its conjunction with the sun, show that the light echo has eroded dramatically. The eastern half is much fainter than in May, and the central hole has extended ∼ 15′′ to the northwest (cf. IAUC 7889). The bright enhancement to the northwest and the ‘fan’ to the southwest (which marked the light-echo appearance in April and May) are, however, still visible. The outer diameter of the light echo is now ≈ 59′′. Magnitudes of the central star (± 0.05 mag due to airmass of 3.5), from observations obtained with USNO 1.0-m and 1.55-m telescopes: Aug. 17.7, J = 6.33, H = 5.42, K′ = 4.37, L′ = 2.77; Aug. 22.50, V = 16.04, R = 14.67, I = 11.17. For comparison, the last measurements before conjunction with the sun gave V = 14.98, R = 12.75, and I = 9.82 on May 16.13 for V838 Mon, and in quiescence it was at V = 15.5.

COMET 46P/WIRTANEN Total visual magnitude estimates: Aug. 13.08 UT, 10.4 (R. J. Bouma,

Groningen, The Netherlands, 0.31-m reflector); 15.11, 10.5 (P. Morel, Eygalayes, France, 0.53-m reflector); 17.89, 10.2 (A. Pearce, Noble Falls, W. Australia, 0.20-m reflector); 20.77, 10.2 (Y. Nagai, Yamanashi, Japan, 0.32m reflector).

COMET C/2000 C7 (SOHO) Contrary to the remarks on IAUC 7919, comet C/2000 C7 is a member

of the Marsden group (cf. IAUC 7832).

2002 August 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green